Circular No. 8962

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 O1 (SOHO) Further to IAUC 8961, a Kreutz sungrazing comet has been found

on SOHO-LASCO coronagraph C3 website images by H. Su (discovery observation tabulated below); according to K. Battams, the object is very small and essentially stellar in appearance with a hint of elongation (mag ∼ 7.5 on July 31.738 UT).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 O1 July 31.596 8^{h}23.^{m}2 +16^{o}01′ C3 HS 2008-O67

This relatively bright Kreutz fragment might be observable during the total solar eclipse on Aug. 1.39-1.47 near α = 8^{h}38^{m}, δ = +16^{o}.8 (equinox 2000.0). The following very preliminary parabolic orbital elements from 16 observations were published on MPEC 2008-P01, together with two ephemerides to indicate the uncertainty in the predicted position during totality.

T = 2008 Aug. 1.68 TT  ω = 91.^{o}48
Ω = 11.90 }2000.0
q = 0.0042 AU  i = 143.38

2007 TY_{430}

S. S. Sheppard, Carnegie Institution of Washington; and C. A. Trujillo, Gemini Observatory, report the discovery of a satellite to the transneptunian minor planet 2007 TY_{430} (cf. MPEC 2008-M38) on images taken on 2007 Oct. 14 at Mauna Kea; images were obtained on several dates with the 8.3-m Subaru, 8.2-m Gemini, and 6.5-m Magellan telescopes. At discovery, the binary had a separation of ∼ 0′′.6 (or at least 13000 km between components) in p.a. ∼ 230^{o}. The two components appear to have a brightness within ∼ 0.1 mag of each other. The individual components’ g′–i′ colors also appear to be very similar. Ongoing monitoring of the binary shows a period of well over 1 year.

THE EDGAR WILSON AWARD 2008 The Smithsonian Astrophysical Observatory announces that the 2008

Edgar Wilson Award for the discovery of comets (cf. IAUC 6936, 8854) is being divided among the following two individuals for the discovery of C/2008 C1: Tao Chen (Suzhou City, Jiangsu province, China) and Xing Gao (Urumqi, Xinjiang province, China).

2008 August 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green